The absolute value of the difference of the squares of numbers that turn up when a pair of fair dice is rolled.

1
Expert's answer
2022-03-03T18:07:29-0500

The largest square is "6^2=36". The smallest square is "1^2=1". So, the largest abosulute difference is "36-1=35"